OpenStudy (anonymous):

Solve : x^4-9x^2+8=0. Thanks.

OpenStudy (aaronq):

x^4-9x^2+8=0 x^4-9x^2=-8 x^2(x^2-9)=-8 x^2[(x-3)(x+3)]=-8 can you take it form here?

OpenStudy (anonymous):

Yes, thanks.

OpenStudy (aaronq):

awesome, no problem.